f(x)=x^2+2x+4 if x>1 or x=1, and f(x)= 3x+4 if x<1 and g(x)=x-3, solve f(g(x)).
Answer by stanbon(75887)

f(x)=x^2+2x+4 if x>1 or x=1, and f(x)= 3x+4 if x<1 and g(x)=x-3, solve f(g(x)).
--------------------
f(g(x)) = f(x-3) = (x-3)^2+2(x-3)+4 = x^2-6x+9 + 2x-6 + 4 = x^2-4x+7
---------------
Cheers,
Stan H.
-----------